Session: Jun 2013 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (C. Nocita at Thu May 23 13:07:07 2013) Success:

Hartebeesthoek (Jonathan Quick at Thu May 23 13:13:53 2013) Success:
No known problems

Medicina (G. Maccaferri at Thu May 23 13:17:23 2013) Success:

Urumqi (Mingfei Gao at Thu May 23 13:28:40 2013) Success:

Svetloe (Shumilova at Thu May 23 16:04:38 2013) Failure:
Electric power failure - only 2 scans recorded

Badary (Podlosinskaya at Thu May 23 16:05:05 2013) Success:

Zelenchukskaya (Ptitcin A. at Thu May 23 16:05:30 2013) Success:

Torun (Kaz Borkowski at Thu May 23 19:04:25 2013) Success:
No known problems with analogue BBC recording. We have also recorded, for the first time, parallelly with Mark5B + DBBC. To our delight there were nice fringes (somewhat higher than from Mark5A), although in only one polarization and in cross-hand correlation. The problem with the other polarization is known and is not related to the new backend itself. I have congratulated Roman Feiler whose efforts made this success possible.

Shanghai (Bo Xia at Fri May 24 08:29:10 2013) Success:

Yebes (Daniel Valero at Fri May 24 08:33:07 2013) Minor failures: TSYS
Wrong computation of Tsys due to an error in preob. Tsys was OK for all the experiment, but the values in the log are not reliable except for the last scan. The error was fixed in preob by the end of the experiment.

Onsala (Michael Lindqvist at Sun May 26 13:15:26 2013) Success:
Used both the DBBC and the old rack. Wrong pol. for the old rack.

Effelsberg (Georgi at Mon May 27 08:00:57 2013) Success:

Westerbork (Geert Kuper at Mon May 27 18:57:12 2013) Success:

Irbene (Vladislavs Bezrukovs at Fri Jul 05 09:05:34 2013) Success:
Receiver LCP channel not working